# size of isogeny classes of abelian varieties in the extended lubin-tate case

### v1

@@ 7. Estimates
The goal of this section is to find the asymptotics for $cl(R_{n})$. By proposition 6.2, this gives a lower bound for $N(A,q^{n}).$ We prove the following estimate.
@@
For a density one set of positive integers, the lower bound for $cl(R_{n})$ is given by $q^{n{(g^{2}-2g/2)}}.$
@@
The class group $\operatorname{cl}(R_{n})$ is approximated the square root of the discriminant of $R_{n}$. We first write down the discriminant of ${\mathbb{Z}}[\alpha].$ Let $\alpha^{\prime}_{i}$, where $1\leq i\leq 2g$ be the conjugates of $\alpha.$ For each $i$, write $\alpha^{\prime}_{i}=q^{1/2}(\cos{\theta}+i\sin{\theta}).$ Then discriminant of ${\mathbb{Z}}[\alpha^{n}]$ is given by
$\prod_{i<j\leq 2g}({\alpha^{\prime}_{i}}^{n}-{\alpha^{\prime}_{j}}^{n})^{2}=q^{\frac{2ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{i})-\cos(n\theta_{j})-i\sin(n\theta_{j}))^{2}$
Next, we estimate the index of ${\mathbb{Z}}[\alpha^{n}]$ in $R_{n}.$ The $\operatorname{disc}R_{n}$ is then estimated by dividing the two estimates. Note that the global index is the product of the local indices. Since the index is $1$ at all primes $l\neq p$, it is sufficient to find a bound for the index after tensoring the orders with ${\mathbb{Z}}_{p}$. We further split the calculation in two parts as follows.
7.0.1. Local analysis.
We make the argument for $n=1$ and the general result follows similarly. Under the embedding of ${\mathbb{Z}}[\alpha]\otimes{\mathbb{Z}}_{p}$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}},$ the image of $\alpha$ is $(\alpha_{1},\alpha_{2})$ where $v_{{\mathfrak{q}}}(\alpha_{1})=\frac{1}{g}$ and $v_{q}(\alpha_{2})=\frac{g-1}{g}$. Locally we are reduced to finding the index of ${\mathbb{Z}}_{p}[(\alpha_{1},\alpha_{2})]$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$. The ring ${\mathbb{Z}}_{p}[(\alpha_{1},\alpha_{2})]$ is monogenic and is isomorphic to ${\mathbb{Z}}_{p}[x]/(f_{1}f_{2})$ where $f=f_{1}f_{2}$ is the factorization of the minimal polynomial $f$ of $\alpha$ in ${\mathbb{Z}}_{p}.$ To find the discriminant of the element, we need to take the difference of the roots of $f_{1}f_{2}$. Let $\beta_{i}$ and $\gamma_{i}$ be the roots of $f_{1}$ and $f_{2}$, that is, the conjugates of $\alpha$. We first divide (the image of) $\alpha$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$ by an appropriate power of $q$ so that the valuation becomes 00. The new element, which we call by $\beta$, gives a monogenic ring ${\mathbb{Z}}_{p}[\beta]$ that is isomorphic to ${\mathbb{Z}}_{p}[x]/g_{1}g_{2}.$ The roots of $g_{1}$ and $g_{2}$ are $\beta_{i}/q^{1/g}$ and $\gamma_{i}/q^{g-1/g}$ for $1\leq i\leq g.$ Thus, we have ${\mathbb{Z}}_{p}[\alpha]\subset{\mathbb{Z}}_{p}[\beta]\subset{\mathcal{O}}_{K}.$ The index $[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]$ is computed using the following relation.
$\operatorname{disc}{\mathbb{Z}}_{p}[\beta]\cdot[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\alpha]$
Up to a unit, the index is given by the following expression.
$\frac{\prod_{1\leq i<j\leq g}(\beta_{i}-\beta_{j})(\gamma_{i}-\gamma_{j})(\beta_{i}-\gamma_{j})}{\prod_{1\leq i<j\leq g}(\beta_{i}/q^{1/g}-\beta_{j}/q^{1/g})(\gamma_{i}/q^{(g-1)/g}-\gamma_{j}/q^{(g-1)/g)})(\beta_{i}/q^{1/g}-\gamma_{j}/q^{g-1/g})}$
$=(q^{1/g})^{g(g-1)/2}\cdot(q^{(g-1)/g})^{g(g-1)/2}\cdot(q^{1/g})^{g^{2}}=q^{\frac{g^{2}+g}{2}}$
We next show that the index ${\mathbb{Z}}_{p}[\beta^{n}]\subset{\mathcal{O}}_{K}$ stays constant for most $n$. We consider ${\mathbb{Z}}_{p}[\beta]\subset{\mathcal{O}}_{K}.$ We need to show that for most $n$, the index of ${\mathbb{Z}}_{p}[\beta]$ is the same as that of ${\mathbb{Z}}_{p}[\beta^{n}].$ We do this by computing the discriminant of $\beta^{n}$ and showing that it’s the same for most $n$. We denote by $\omega_{i}$ the conjugates of $\beta$, that is the roots of $g_{1}$ and $g_{2}$ as described above.
@@
$\operatorname{disc}{\mathbb{Z}}_{p}[\beta^{n}]=\prod_{1\leq i<j\leq 2g}(\omega_{i}^{n}-\omega_{j}^{n})^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\beta]\cdot\prod_{1\leq i<j\leq 2g}(\omega_{i}^{n-1}+\omega_{i}^{n-2}\omega_{j}+\dots+\omega_{j}^{n-1})$
Now, note that the residue field of ${\mathcal{O}}_{K_{1}}$ (and ${\mathcal{O}}_{K_{2}}$) is ${\mathbb{F}}_{p^{g}}.$ So if we take $n$ larger than $p^{g}-1$ then all the conjugates of $\beta$ are $1\operatorname{mod}$ $p$. Moreover, if we take $n$ co-prime to $p$, then the discriminant of ${\mathbb{Z}}_{p}[\beta^{n}]$ and ${\mathbb{Z}}_{p}[\beta]$ differ by a unit. This proves that the index of ${\mathbb{Z}}[\beta^{n}]\subset R_{n}$ stays the same for a density one set of positive integers $n$.
7.0.2. Back to the global setting.
Up to some constant, the discriminant of $R_{n}$ is bounded above by the following expression.
$\frac{q^{\frac{ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{j})-\cos(n\theta_{j})-\sin(n\theta_{j}))^{2}}{q^{n(g^{2}+g)}}$
@@
Since the term with sines and cosines is bounded, we see that the class group $cl(R_{n})$ should be approximated by $q^{\frac{2ng(2g-1)}{2}-\frac{n(g^{2}+g)}{2}}=q^{n(g^{2}-2g)/2}$
@@
For a positive density set of positive integers $n$, the size of (unpolarised) isogeny class of $\widetilde{A}$ in characteristic zero is greater than or equal to $q^{n(\frac{(g^{2}-2g)}{2}+o(1))}$.
@@ 8. Polarisation
So far we have computed a lower bound for the number of isomorphism classes of simple abelian varieties with endomorphism ring $R_{n}$ and CM-type $\Phi.$ In order to restrict this count to the isogeny class in $\mathcal{A}_{g,1},$ we need to take polarisation into account. Howe in [How95] describes the polarisation data for ordinary abelian varieties which is summed up nicely in [ST18], section 3.1. However, since the data is described purely in characteristic zero and is independent of the type of abelian variety, we use the same description in our case.
@@
We give a brief summary of the description of polarisation on . We fix a CM type $\Phi$ on $K$. The polarisation on an abelian variety corresponding to $I$ in $cl(R_{n})$ is given by a $K$-bilinear form on $K$
$\Psi:(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(\lambda x\overline{y})$
with the following properties to keep in mind.
@@
    Here, $\lambda\in K$ is purely imaginary.
@@
    For all $\varphi\in\Phi$, we have $\varphi(\lambda)/i$ is positive.
@@
We note that if an ideal $I$ is dual to itself with respect to the above form then then its corresponding abelian variety will be principally polarized. To restrict our count to principally polarised abelian varieties, we have the following proposition. The proof of this proposition goes through analogously to Lemma 3.8 in [ST18].
@@
is either empty or is in one-to-one correspondence with principally polarised abelian varieties isogenous to $A$ over ${\mathbb{F}}_{q^{n}}.$ Here $cl^{+}(R_{n}^{+})$ denotes the narrow class group of the order $R_{n}^{+}$ and the norm map refers to $I\mapsto I\overline{I}$.
@@
For a positive density set of integers $n,$ we have $\frac{|cl(R_{n})|}{|cl^{+}(R_{n}^{+})|}=q^{\frac{n(g+1)(g-2)}{2}}$
@@
It suffices to find an upper bound for $cl(R_{n}^{+})$. Using a similar strategy from the previous computation, we will bound the index $[{\mathbb{Z}}[\alpha_{n}^{+}]:R_{n}^{+}].$ For simplicity we do the computation with $n=1.$ Let $\alpha^{+}$ be the element $\pi+q/\pi$. We denote the $p$-adic conjugates of $\alpha^{+}$ by $\alpha_{i}^{+}$ for $1\leq i\leq g.$ Following the same strategy as the previous section, the index of ${\mathbb{Z}}[\alpha^{+}]$ inside $R_{n}^{+}$ estimated by their index at $p$. The index is given by the following expression at $p$.
$\frac{\prod_{1\leq i<j\leq g}(\alpha_{i}^{+}-\alpha_{j}^{+})}{\prod_{1\leq i<j\leq g}(\alpha_{i}/q^{1/g}-\alpha_{j}/q^{1/g})}=q^{\frac{{g-1}}{2}}$
Dividing the discriminant of ${\mathbb{Z}}[\alpha^{+}]$ calculated as in [ST18] with the index obtained above, we get that $cl(R^{+})$ is approximated by $q^{\frac{g^{2}-3g+2}{4}}$. Combining this with the estimate for $cl(R_{n})$ we get the desired result. ∎
We note that if $\iota:\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ is a polarisation on $\widetilde{A}$, then it extends to a polarisation on its Néron model $\mathscr{A}$ and hence, reduces to a polarisation on the special fibre as well. Therefore, the count in characteristic zero restricted to principally polarised varieties gives us the right estimate in characteristic $p$. It remains to prove that the kernel of the Norm map in Proposition 8.1 is non empty. That is, the subset of $I(\widetilde{A},L^{\prime})$ with endomorphism ring $R_{n}$ is not empty. This is proved in the final proposition of this paper.
@@
We assume that $q^{1/2},\alpha=\alpha_{1},$ and its conjugates $\alpha_{2},\dots\alpha_{g}$, are multiplicatively independent. Then there exist principally polarised abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ for a positive density set of integers $n.$
@@
Let $R^{\prime}_{n}={\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and let $h_{n}$ denote the minimal polynomial of $\alpha^{n}+(q/\alpha)^{n}$ and let $\lambda_{n}=(\alpha^{n}-(q/\alpha)^{n}h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n}))^{-1}.$ By Proposition 9.4 in [How95], the trace dual of $R^{\prime}$ is $\lambda_{n}^{-1}R^{\prime}.$ Therefore, if we consider the bilinear form
$(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(\lambda_{n}x\overline{y})$
@@
then the dual of $R^{\prime}_{n}$ with respect to the form is itself. Furthermore, the above form induces a polarisation on the elements of $cl(R_{n})$ when $\varphi(\lambda_{n})/i$ is a positive real number for every $\varphi\in\Phi.$ The argument given in Proposition $3.6$ of [ST18] shows that for a positive proportion of $n$, $\varphi(\lambda_{n})/i$ is positive under the assumption that the set $\{q^{1/2},\alpha_{1},\dots,\alpha_{g}\}$ is multiplicatively independent. We assume that $n$ is an integer such that $\lambda_{n}$ satisfies appropriate conditions so that we get a polarisation on the elements of $cl(R_{n}).$ Let ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}$ be prime ideals in ${\mathcal{O}}_{K}$ lying over $p.$ Without loss of any generality, we assume $n=1$. Since $R=R_{1}$ is maximal at ${\mathfrak{p}}_{1}$, the ideal $R\cap{\mathfrak{p}}_{1}$ is invertible. The same holds for ${\mathfrak{p}}_{2}$. By abuse of notation we will use the same notation for the $R$-ideals $R\cap{\mathfrak{p}}_{1}$ and $R\cap{\mathfrak{p}}_{2}.$ In the argument that follows, we will go back and forth between abelian varieties and their corresponding lattice in $K$ using Proposition 6.2. We know that the lattice $R_{n}$ of $\widetilde{A}$ over ${\mathbb{Z}}_{p}$ admits a decomposition corresponding to the slope $1/g$ and $g-1/g$ of (the special fibre of) $\widetilde{A}.$ We write this decomposition as $T_{1/g}\oplus T_{g-1/g}.$ We recall that $T_{1/g}$ (and $T_{g-1/g}$) has endomorphism by the maximal order in $K_{1}$ (and $K_{2}$). Now, $R^{\prime}$ and $R$ agree everywhere except at ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2},$ and hence the same is true for $R^{\prime}$ and the dual of $R$. Therefore, the dual of $R$ is ${\mathfrak{p}}_{1}^{m}{\mathfrak{p}}_{2}^{k}$ for some integers $m$ and $k.$ The kernel of the corresponding polarisation map $\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ corresponds to the $\overline{\omega}^{m}$-torsion of $T_{1/g}$ and $\overline{\omega}^{k}$-torsion of $T_{g-1/g}$ where $\overline{\omega}$ is the uniformizer of ${\mathcal{O}}_{K_{1}}$ and ${\mathcal{O}}_{K_{2}}$. Since the polarisation map has a kernel that is self dual, we see that $m=k.$ Therefore, the kernel equals to the $p^{gm}$-torsion of $\widetilde{A}$ and hence $\widetilde{A}$ is principally polarised as $\widetilde{A}^{\vee}\simeq\widetilde{A}/\widetilde{A}[p^{gm}]\simeq\widetilde{A}$.

### v2

@@ 6. Estimates
The goal of this section is to find the asymptotics for $Cl(R_{n})$. We prove the following estimate.
@@
For a positive density set of positive integers $n$, the class group $Cl(R_{n})=q^{n{((g^{2}-2g)/2+o(1))}}.$
@@
As $n$ grows large, the class group of an order is well approximated by the square root of its discriminant. Let $\alpha_{i}$, where $1\leq i\leq 2g$ and $\alpha_{1}=\alpha$ be the conjugates of $\alpha.$ For each $i$, we write $\alpha_{i}=q^{1/2}(\cos{\theta}+i\sin{\theta}).$ Then discriminant of the monogenic order ${\mathbb{Z}}[\alpha^{n}]$ is given by
$\operatorname{disc}({\mathbb{Z}}[\alpha^{n}])=\prod_{1\leq i<j\leq 2g}({\alpha^{n}_{i}}-{\alpha^{n}_{j}})^{2}=q^{\frac{2ng(2g-1)}{2}}\prod_{1\leq i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{i})-\cos(n\theta_{j})-i\sin(n\theta_{j}))^{2}$
Since $\operatorname{disc}({\mathbb{Z}}[\alpha^{n}])=[{\mathbb{Z}}[\alpha^{n}]:R_{n}]^{2}\operatorname{disc}(R_{n})$, it is enough to estimate the index of ${\mathbb{Z}}[\alpha^{n}]$ in $R_{n}.$ We note that the index is $1$ at all primes $l\neq p$ (see Definition 3.3, point (3)), hence it is sufficient to estimate the index after tensoring the orders with ${\mathbb{Z}}_{p}$. This is done below.
6.1. Estimating the index at the prime $p$
We write ${\mathbb{Z}}[\alpha]\otimes{\mathbb{Z}}_{p}\subset{\mathcal{O}}_{K}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. Under this embedding, the image of $\alpha$ is $(\beta,\gamma)$ where $v_{{\mathfrak{q}}}(\beta)=1/g$ and $v_{q}(\gamma)=(g-1)/g$. The ring ${\mathbb{Z}}_{p}[(\beta,\gamma)]$ is monogenic and is isomorphic to ${\mathbb{Z}}_{p}[x]/(f_{1}f_{2})$ where $f=f_{1}f_{2}$ is the factorization of the minimal polynomial $f$ of $\alpha$ in ${\mathbb{Z}}_{p}.$ Let $\beta_{i}$ and $\gamma_{i}$ be the roots of $f_{1}$ and $f_{2}$ with $\beta_{1}=\beta$ and $\gamma_{1}=\gamma$ and $1\leq i\leq g$. Now, for any positive integer $n$, we are reduced to finding the index of ${\mathbb{Z}}_{p}[\alpha^{n}]={\mathbb{Z}}_{p}[(\beta^{n},\gamma^{n})]$ in ${\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. We compute this as follows:
Consider the monogenic ring ${\mathbb{Z}}_{p}[x]/g_{1}g_{2}:={\mathbb{Z}}_{p}[\delta]$ where $\delta=(\beta/q^{1/g},\gamma/q^{g-1/g})$ and the roots of $g_{1}$ and $g_{2}$ are given by $\beta^{\prime}_{i}:=\beta_{i}/q^{1/g}$ and $\gamma^{\prime}_{i}:=\gamma_{i}/q^{g-1/g}$ respectively where $1\leq i\leq g$. Thus, we have the following containment of orders:
${\mathbb{Z}}_{p}[\alpha^{n}]\subseteq{\mathbb{Z}}_{p}[\delta^{n}]\subseteq{\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$
Now for most $n$, the index ${\mathbb{Z}}_{p}[\delta^{n}]\subseteq{\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ remains constant, that is does not depend on $n$. Indeed, the discriminant of ${\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ is constant, while that of ${\mathbb{Z}}_{p}[\delta^{n}]$ is
@@ 6.1. Estimating the index at the prime $p$
$\operatorname{disc}({\mathbb{Z}}_{p}[\delta^{n}])=\prod_{1\leq i<j\leq g}((\beta^{\prime}_{i})^{n}-(\beta^{\prime}_{j})^{n})((\gamma^{\prime}_{i})^{n}-(\gamma^{\prime}_{j})^{n})\prod_{1\leq i,j\leq g}((\beta^{\prime}_{i})^{n}-(\gamma^{\prime}_{j}))$
$={\mathbb{Z}}_{p}[\delta]\prod_{1\leq i<j\leq g}((\beta^{\prime}_{i})^{n-1}+(\beta^{\prime}_{i})^{n-2}\beta^{\prime}_{j}+\dots+(\beta^{\prime}_{1})^{n-1})((\gamma^{\prime}_{i})^{n-1}+(\gamma^{\prime}_{i})^{n-2}\gamma^{\prime}_{j}+\dots+(\gamma^{\prime}_{1})^{n-1})$
$((\gamma^{\prime}_{i})^{n-1}+(\gamma^{\prime}_{i})^{n-2}\beta^{\prime}_{j}+\dots+(\beta^{\prime}_{j})^{n-1})$
We recall that $K^{\prime}$ equals the unramified extension of ${\mathbb{Q}}_{p}$ of degree $g$ and hence, for $n$ divisible by $p^{g}-1$ then all $\beta_{i}$ and $\gamma_{j}$ with $1\leq i,j\leq g$ are congruent to $1\operatorname{mod}$ $p$. Moreover, if we take $n$ co-prime to $p$, then the discriminant of ${\mathbb{Z}}_{p}[\beta^{n}]$ and ${\mathbb{Z}}_{p}[\beta]$ differ by a unit.
Hence, for such $n$ it is enough to compute the index of ${\mathbb{Z}}[\alpha^{n}]$ in ${\mathbb{Z}}[\delta^{n}]$. This is is straightforward using the relation:
$\operatorname{disc}{\mathbb{Z}}_{p}[\delta^{n}]\cdot[{\mathbb{Z}}_{p}[\alpha^{n}]:{\mathbb{Z}}_{p}[\delta^{n}]]^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\alpha^{n}]$
The index is given by the following expression for $n=1$ and can be raised to the power $n$ for any positive integer $n$.
$[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]=\frac{\prod_{1\leq i<j\leq g}(\beta_{i}-\beta_{j})(\gamma_{i}-\gamma_{j})\prod_{1\leq i,j\leq g}(\beta_{i}-\gamma_{j})}{\prod_{1\leq i<j\leq g}(\beta_{i}/q^{1/g}-\beta_{j}/q^{1/g})(\gamma_{i}/q^{(g-1)/g}-\gamma_{j}/q^{(g-1)/g)})\prod_{1\leq i,j\leq g}(\beta_{i}/q^{1/g}-\gamma_{j}/q^{(g-1)/g})}$
which equals $(q^{1/g})^{g(g-1)/2}\cdot(q^{(g-1)/g})^{g(g-1)/2}\cdot(q^{1/g})^{g^{2}}=q^{(g^{2}+g)/2}$ up to some constant. Hence, for a positive density set of positive integers, we get $[{\mathbb{Z}}[\alpha^{n}]:R_{n}]=[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]=q^{n(g^{2}+g)/2}$ up to some constant.
6.2. Final estimate
For a positive density set of positive integers $n$, the discriminant of $R_{n}$ equals (up to some constant)
$\frac{(\operatorname{disc}({\mathbb{Z}}[\alpha^{n}]))^{1/2}}{[{\mathbb{Z}}[\alpha^{n}]:R_{n}]}=\frac{q^{\frac{ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{j})-\cos(n\theta_{j})-\sin(n\theta_{j}))^{2}}{q^{n(g^{2}+g)/2}}$
@@ 6.2. Final estimate
Since the term with sines and cosines is bounded, we see that the class group $Cl(R_{n})$ is well approximated by the leading term $q^{\frac{ng(2g-1)}{2}-\frac{n(g^{2}+g)}{2}}=q^{n(g^{2}-2g)/2}$. ∎
@@
For a positive density set of positive integers $n$, the size of isogeny class of $A$ (without any conditions on polarisation) in characteristic zero is greater than or equal to $q^{n(\frac{(g^{2}-2g)}{2}+o(1))}$.
Proof.
This follows from Corollary 5.3 and the calculation above in Proposition 6.1. ∎
@@ 7. Polarisation
So far, we have an estimate for the size of the isogeny class without specifying any condition on the polarisation. In order to restrict this count to the isogeny class in $\mathcal{A}_{g,1},$ we need to impose further conditions for principal polarisation. In [How95], Howe describes the polarisation data for ordinary abelian varieties which is summed up nicely in [ST18], Section 3.1. However, since the data is described purely in characteristic zero and is independent of the type of abelian variety, we use the same description in our case that we describe below.
@@
We work with the CM type $\Phi$ on $K$ as before. For any element $x\in K$, we denote $\overline{x}$ to be the complex conjugation (induced by $K$) on $x$. The polarisation on an abelian variety corresponding to the ideal $I$ in $Cl(R_{n})$ (see Proposition 5.2) is given by a ${\mathbb{Q}}$-valued bilinear form on $K$.
$\Psi:(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(ux\overline{y})$
satisfying the following properties.
@@
    The form $\Psi$ restricted to $I$ is valued in ${\mathbb{Z}}$.
@@
    The element $u\in K$ is purely imaginary.
  3. (3)
    For all $\varphi\in\Phi$, we have $\varphi(u)/i$ is positive.
@@
We note that if an ideal $I$ is dual to itself with respect to the above form then then its corresponding abelian variety will be principally polarized. We have the following proposition.
@@
is either empty or is in one-to-one correspondence with principally polarised abelian varieties isogenous to $\widetilde{A}$ over $L^{\prime}$ with endomorphism ring $R_{n}$ (and hence isogenous to $A$ over ${\mathbb{F}}_{q^{n}}$). Here $Cl^{+}(R_{n}^{+})$ denotes the narrow class group of the order $R_{n}^{+}$ and the norm map refers to $I\mapsto I\overline{I}$.
Proof.
The proof of this proposition goes through verbatum as in Lemma 3.8 in [ST18]. ∎
@@
For a positive density set of integers $n,$ we have $\frac{|Cl(R_{n})|}{|Cl^{+}(R_{n}^{+})|}=q^{(n(g+1)(g-2)/2+o(1))}$
@@
It suffices to find an upper bound for $Cl^{+}(R_{n}^{+})$. Let $\lambda^{(n)}$ be the element $\alpha^{n}+q^{n}/\alpha^{n}$. We denote the $p$-adic conjugates of $\lambda^{(n)}$ by $\lambda^{(n)}_{i}$ for $1\leq i\leq g$ and $\lambda^{(n)}_{1}:=\lambda^{(1)}$. As before, in order to estimate the class group, we first estimate the index $[{\mathbb{Z}}[\lambda^{(n)}]:R_{n}^{+}].$ For simplicity we do the computation with $n=1$ as for a general $n$, the estimate is raised to the power of $n$. We write $\lambda^{(1)}_{i}=\lambda_{i}$ for $1\leq i.j\leq g$ and $R^{+}_{1}:=R^{+}$. It is enough to compute the index of ${\mathbb{Z}}[\lambda]$ in $R^{+}$ at the prime $p$. We recall $K^{\prime}$ is the unramfied extension of ${\mathbb{Q}}_{p}$ of degree $g$. As $R^{+}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}$, we can follow the the same argument as in the proof of Proposition 6.1 to compute this. Hence, for a positive proportion of positive integers $n$ (satisfying the same congruent conditions as in the proof of Proposition 6.1), the index is given by the following expression:
$[{\mathbb{Z}}[\lambda]:R^{+}]=\frac{\prod_{1\leq i<j\leq g}\lambda_{i}-\lambda_{j}}{\prod_{1\leq i<j\leq g}(\lambda_{i}/q^{1/g}-\lambda_{j}/q^{1/g})}=q^{\frac{{g-1}}{2}}$
The discriminant of the monogenic ring ${\mathbb{Z}}[\lambda]$ is calculated easily, and is also given in Lemma 3.8 in [ST18] where the leading the term is given by $q^{g(g-1)/2}.$ Now, the leading term of the $\operatorname{disc}(R^{+})$ is approximated by dividing this discriminant by the square of the index computed above. Hence $Cl(R^{+})$ comes out to be $q^{(g^{2}-3g+2)/4}$, that is taking the square root of $\operatorname{disc}(R^{+})$. Combining this (raised to the power of $n$) with the estimate of $Cl(R_{n})$ in Proposition 6.1 we get the desired result. ∎
We note that a polarisation on $\widetilde{A}$ extends to a polarisation on the Néron model and hence, reduces to a polarisation on its special fibre $A$ as well. Therefore, the count in characteristic zero restricted to principally polarised varieties gives us the right estimate in characteristic $p$. It remains to prove that the kernel of the Norm map in Proposition 7.1 is non empty. That is, the subset of $I(\widetilde{A},L^{\prime})$ with endomorphism ring $R_{n}$ is not empty. This is proved in the final proposition of this paper.
@@
We assume that $q^{1/2},\alpha=\alpha_{1},$ and its conjugates $\alpha_{2},\dots\alpha_{g}$, are multiplicatively independent. Then for a positive density set of integers $n$, there exist principally polarised abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$.
@@
Let $h_{n}$ denote the minimal polynomial of $\alpha^{n}+(q/\alpha)^{n}$ and $h_{n}^{\prime}$ be its derivative. Let $u_{n}=(\alpha^{n}-(q/\alpha)^{n}h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n}))^{-1}$. We note that $\alpha^{n}-(q/\alpha)^{n}$ is purely imaginary while $h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n})$ is purely real. Hence $u_{n}$ is purely imaginary for all $n$. By (the proof of) Proposition 9.4 in [How95], the trace dual of ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ is $u_{n}^{-1}{\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$. Therefore, if we consider the bilinear form
$(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(u_{n}x\overline{y})$
@@
then the dual of ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ with respect to this form is itself. Furthermore, the above form induces a polarisation on the elements of $Cl(R_{n})$ when $\varphi(u_{n})/i$ is a positive real number for every $\varphi\in\Phi.$ The argument given in Proposition $3.6$ of [ST18] shows that for a positive proportion of $n$, $\varphi(u_{n})/i$ is positive under the assumption that the set $\{q^{1/2},\alpha_{1},\dots,\alpha_{g}\}$ is multiplicatively independent. We assume that $n$ is an integer such that $u_{n}$ satisfies appropriate conditions so that we get a polarisation on the elements of $Cl(R_{n}).$ Without loss of any generality, we assume $n=1$ as the argument remains the same for any $n\geq 1$. Let $\widetilde{A}/L^{\prime}$ be the abelian variety corresponding to the order $R_{1}:=R$ and let $A$ be its reduction mod $p$ over ${\mathbb{F}}_{q}$. Further, let ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}$ be the intersection of prime ideals in ${\mathcal{O}}_{K}$ lying over $p$ with $R$. Since $R$ is locally maximal at the primes lying over $p$, the prime ideals ${\mathfrak{p}}_{1}$ and $\mathfrak{p}_{2}$ are invertible. Note that ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and $R$ agree everywhere except at ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2},$ and hence the same is true for ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]^{\vee}={\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and $R^{\vee}$. Therefore, $R$ is ${\mathfrak{p}}_{1}^{m}{\mathfrak{p}}_{2}^{k}$, while $R^{\vee}={\mathfrak{p}}_{1}^{k}{\mathfrak{p}}_{2}^{m}$ for some integers $m$ and $k.$ Now, the lattice $R$ locally admits a decomposition corresponding to the slope $1/g$ and $g-1/g$ of $A$. We write this decomposition as $T_{1/g}\oplus T_{g-1/g}.$ We recall that $T_{1/g}$ and $T_{g-1/g}$ have endomorphism by the maximal order in $K^{\prime}$ where $K^{\prime}$ is the unramified extension of ${\mathbb{Q}}_{p}$ of degree $g$. Hence, the kernel of the corresponding polarisation map $\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ corresponds to the direct sum of $\overline{\omega}^{m}$-torsion of $T_{1/g}$ and $\overline{\omega}^{k}$-torsion of $T_{g-1/g}$ where $\overline{\omega}$ is the uniformizer of ${\mathcal{O}}_{K^{\prime}}$. Since the polarisation map has a kernel that is self dual, we see that $m=k.$ Therefore, the kernel equals to the $p^{gm}$-torsion of $\widetilde{A}$ and hence $\widetilde{A}$ is principally polarised as $\widetilde{A}^{\vee}\simeq\widetilde{A}/\widetilde{A}[p^{gm}]\simeq\widetilde{A}$. ∎